Included in Software/LBCControlSoftwareNotes page

LBC Software Issue Trak Status

IT Detail Status
7032 Red HK not coming up (timeout talking to dataprobe) the iBootBar dataprobes have sometimes required someone to log in and type reboot and we haven't figured out what gets it into this state.
One thing to do is to make sure we are cleaning up the socket to the device everywhere, even on errors.
6800 OB hung with R (8.0s guiding failed) and B (GOYA rotator warning) no issue with the rotator, this problem is red controller had a hardware failure but somehow missed the STOP command that the software issued to recover. The observer issued a STOP command, but didn't wait until it was done to issue another STOP.
The hang should be fixed in version 2.15, released Sep-2018.
6756 LBC pause/resume behavior UI reported number of exposures taken as greater than total requested
why is this only showing up when using one channel?
5715 stars trailed on first LBC image after a dither John says this is likely due to LBC not checking for collimation update to finish; this is a one-sided LBC problem
Should be fixed in version 2.15, released Sep-2018.
5302 Suggested changes/reorg of LBC image headers enhancement
5205 OB progress is incorrect after pause/resume bug
5065
5548
Timeout error on motion of filter wheel We see this with wheel #2 for both red and blue. Seems like mostly it's a move command that doesn't get received on the motor because we timeout and then check and it hasn't moved at all.
4705 Add pointing offsets to LBC headers enhancement
Waiting on Olga for specifics, then CB can add to the TCS public interface.
4642 Add notification on failures enhancement
pop up on tracking stopped - no, we do NOT want to stop operations
added ERROR message when tracking background is < 1 should it be identically zero? - cannot do a pop-up without stopping the OB. Fixed so it waits til it has tried up to the max exposure time before it complains
Added Z4 to GUI.
Make the background of the ? red -- needs testing.
Olga also added the "shutter seems to be not working" to this IT - changed this to return an ERROR - so it does not go to Ready for the camera anymore -- tested it.
how do we test the science image? don't have to, they don't have this problem
4566 Red Rotator communication problem bug : When the rotator failed in a preset, it tried to retry indefinitely - it re-tried ~680 times before they stopped it. It is not really obvious to the operator that is happening, because it automatically retries and claims rotator system recovered from a problem ; The one error they see is: >>> preset/expose/save interrupted at preset time
also core-dumped on the first lbcstart they did after this problem (this is not uncommon, often core's on startup) at least the startup core is fixed!
4236 hung at unitialization of blue HK this is the HK timeouts - cannot seem to fix this within the software, but a lbckill/lbcstart fixes it!
3833 audible alarm on errors enhancement
3816 Scale R/B exposures separately enhancement
3759 Blue tracker images suddenly all zeroes bug
Happens on both sides. Goes away with an OFF/ON toggle. Sometimes it's not zero, but just funky.
I have added an ERROR in the log when the background goes to zero check the fits file for zero mean (releasing Feb-2016); added a failed state to tracker status so the GUI can turn the "?" red. need to test it
Also look at adding a "reset" that power-cycles the electronics, but not the windows PCs.
3595 Rotator position on first dither still seeing problems as of Jan-2012
2765 tick marks are sticky on the UI I did some testing - this is all controlled by the browser settings and cookies and what you submit to the GUI. If you check mark, but do not "submit" by pressing PLAY or TURN ON or TURN OFF, then the settings will reset to the last time that happened in this browser environment. If you clear and restart, nothing is sticky - if you kill and restart, it remembers.
Olga notes in the IT that running dofpia and the check marks erasing? DOFPIA "submits" the http form so the checkmarks will fill in based on that.
2631 too few downloaded pixels error This is not mine, but I am monitoring the problem.
2573 LBCB/LBCR First Image of a Sequence has small Rotator Offset (~0.5 deg.) reopened in Sep-2013; AdP says maybe to track Parallactic angle to open shutter delay.
2546 WCS info in LBC tech chip headers enhancement
Waiting on Olga to provide
2397 RPC Client/Host Comm Error bugs
multiple problems documented here - guiding lapses, missing images, inhibit counter on restart, etc.
This became a big issue coming out of summer shutdown 2013. Added better logging, retry during preset.
Need to: release go.bat script that logs when restarting the RPC servers; redirect Windows side output; check all timeouts are long enough; send DD entries in separate thread?
Topic revision: r48 - 11 Sep 2018, KelleeSummers
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback